A=A0e^-0.5t is the equation that gives the decay of a radioactive substance. the amount of the substance at the beginning is A0, and the amount after t years is A. Find the half-life (time for half of the substance to decay) correct to the nearest thousandth.
Answer by nerdybill(7384)

A=A0e^-0.5t is the equation that gives the decay of a radioactive substance. the amount of the substance at the beginning is A0, and the amount after t years is A. Find the half-life (time for half of the substance to decay) correct to the nearest thousandth.
.5A0 = A0e^(-0.5t)
.5 = e^(-0.5t)
ln(.5) = -.0.5t
ln(.5)/(-.0.5) = t
(-0.69314718056)/(-.0.5) = t
1.38629436112 = t
1.386 years = t